An orcharder spends $90,000 to plant pomegranate bushes. It will take four years for the bushes to provide a usable crop. He estimates that every year for 20 years after that he will receive a crop worth $9,500 per year. If the discount rate is 8%​, what is the net present value​ (NPV) of this​ investment?
Answer #1

NPV calculation is as follows:

  • The first inflow comes at the end of year 4 and the last at the end of year 23 making it 20 inflows
Year CF Discount Factor Discounted CF
0 $   -90,000.00 1/(1+0.08)^0= 1 1*-90000=       -90,000.00
1 $                    -   1/(1+0.08)^1= 0.925925926 0.925925925925926*0=                        -  
2 $                    -   1/(1+0.08)^2= 0.85733882 0.857338820301783*0=                        -  
3 $                    -   1/(1+0.08)^3= 0.793832241 0.79383224102017*0=                        -  
4 $       9,500.00 1/(1+0.08)^4= 0.735029853 0.735029852796453*9500=           6,982.78
5 $       9,500.00 1/(1+0.08)^5= 0.680583197 0.680583197033753*9500=           6,465.54
6 $       9,500.00 1/(1+0.08)^6= 0.630169627 0.630169626883105*9500=           5,986.61
7 $       9,500.00 1/(1+0.08)^7= 0.583490395 0.583490395262134*9500=           5,543.16
8 $       9,500.00 1/(1+0.08)^8= 0.540268885 0.540268884501976*9500=           5,132.55
9 $       9,500.00 1/(1+0.08)^9= 0.500248967 0.500248967131459*9500=           4,752.37
10 $       9,500.00 1/(1+0.08)^10= 0.463193488 0.463193488084684*9500=           4,400.34
11 $       9,500.00 1/(1+0.08)^11= 0.428882859 0.428882859337671*9500=           4,074.39
12 $       9,500.00 1/(1+0.08)^12= 0.397113759 0.397113758645991*9500=           3,772.58
13 $       9,500.00 1/(1+0.08)^13= 0.367697925 0.367697924672214*9500=           3,493.13
14 $       9,500.00 1/(1+0.08)^14= 0.340461041 0.340461041363161*9500=           3,234.38
15 $       9,500.00 1/(1+0.08)^15= 0.315241705 0.31524170496589*9500=           2,994.80
16 $       9,500.00 1/(1+0.08)^16= 0.291890468 0.291890467561009*9500=           2,772.96
17 $       9,500.00 1/(1+0.08)^17= 0.270268951 0.270268951445379*9500=           2,567.56
18 $       9,500.00 1/(1+0.08)^18= 0.250249029 0.250249029116092*9500=           2,377.37
19 $       9,500.00 1/(1+0.08)^19= 0.231712064 0.231712063996381*9500=           2,201.26
20 $       9,500.00 1/(1+0.08)^20= 0.214548207 0.214548207404057*9500=           2,038.21
21 $       9,500.00 1/(1+0.08)^21= 0.198655748 0.198655747596349*9500=           1,887.23
22 $       9,500.00 1/(1+0.08)^22= 0.183940507 0.183940507033656*9500=           1,747.43
23 $       9,500.00 1/(1+0.08)^23= 0.170315284 0.170315284290422*9500=           1,618.00
NPV = Sum of all Discounted CF       -15,957.36

As NPV is negative, the project should not be invested in.